on a road map with a scale of 1/4 inch per 10 miles, the highway from Waukee to Winterset is 1 3/8 inches long. How may miles long is this highway?
thank you very much for helping me i have tried but i have forgoten how to work out this problem.

Answer by AnlytcPhil(1116)

on a road map with a scale of 1/4 inch per 10 miles, 
the highway from Waukee to Winterset is 1 3/8 inches 
long. How may miles long is this highway? thank you 
very much for helping me i have tried but i have 
forgoten how to work out this problem.

(1/4 in) is to (10 miles) as (1 3/8 in) is to (x miles)

Replace the "is to"'s by " / " and "as" by " = "

(1/4 in)/(10 miles) = (1 3/8 in)/(x miles)

Erase the units:

         (1/4)/(10) = (1 3/8)/(x)

Cross multiply

           (1/4)(x) = (10)(1 3/8)

Change 1 3/8 to  11/8

           (1/4)(x) = (10)(11/8)

                x/4 = 55/4

                  x = 55

Answer 55 miles

Edwin